# url]https://www.bangalab.org

### v1

@@ 3.1.3 Constraints and the feasible set $\mathcal{Y}$
While the first two types are typically addressed by a modeling expert, the third type is something that needs to be considered in the inverse optimization problem. It comprises vectors of general inequality constraints $g(y)\geq 0$. They allow to formulate lower and upper bounds on variables, but also more general relationships. Of particular importance is whether an inequality constraint $g_{i}(y)\geq 0$ is active in an optimal point, i.e., $g_{i}(y^{*})=0$, or inactive, i.e., $g_{i}(y^{*})>0$. Only active constraints have a local influence on the optimum (i.e., without this constraint the optimal point $y^{*}$ could be different). In a practical setting, if it is unclear wether the data $\eta$ was produced by an optimal, but constrained process, one could simply solve two problems – one with the constraint $g_{i}(y)\geq 0$ on the inner level, and one without. The better match to the data $\eta$ is then the more likely explanation. However, the number of possible combinations grows quickly if there are several candidate constraint functions; therefore, we model the inclusion of constraints again as a degree of freedom for the optimization on the outer level of (1). Assume that $n_{g}$ constraint function candidates $g_{i}(\cdot)$ are known, but it is unknown and of interest, if they were active and played a role in the observed process. One possibility to model them is via binary weight variables $w_{i}\in\{0,1\}$ as
@@ 3.1.5 Further considerations
Optimality principles in biology seem to be nested in a hierarchical way, as discussed above. Contributing to the main goal of maximizing fitness, biological systems may have modi operandi that are optimal for a specific function or task and the considered scale and environmental setting. Let us assume the simplified situation of a lion who knows at least three different kinds of behavior: relaxing (energy consumption minimized), sneaking up on prey (observability minimized), and hunting (time-to-target minimized). Obviously, all of them contribute to the main goal of maximizing fitness. Yet, determining the exact objective function from data may be more difficult if several (switching) objectives were used in the observed time horizon. Such switches can be further classified. They may either be due to external circumstances (for example, the predator suddenly observes a prey and switches from relaxing to foraging mode) or because several consecutive stages like sneaking and hunting are optimized together when foraging. Mathematically, this difference is important. In the first case, two different optimal policies are concatenated, while in the second case the optimal policy considers all stages in one go. The difference will be illustrated in Section 3.2. Additionally, the current situation (modeled via parameters $p$) has an influence. A resting lion will react differently to a prey depending on if it is hungry or not. But in both cases, the overall cost function (fitness towards survival and reproduction) is the same.
@@
In most situations where optimality principles are used to explain (or predict) dynamics, it is sufficient to consider the open loop optimal control problem, where in addition to the cost function we obtain the optimal control policies as a function of time, $u(t)$. For example, in the case of the simple linear metabolic pathway considered by Klipp et al. [55], we can obtain the time-dependent enzyme profiles and the optimality principle consistent with the data by using inverse optimal control [157]. This can be satisfactory if we simply want to explain or predict the observed dynamic behavior. However, if we want to obtain more information about the regulatory mechanisms involved, it should be noted that in reality, the enzyme concentrations depend on the expression of genes, which are themselves influenced (activated or inhibited) by the concentrations of other metabolites and enzymes, or even the expression of other genes. Thus, one might want to infer the regulatory network compatible with the observed behavior. This would be particularly useful in biotechnological and biomedical applications where we seek to change such regulations using genetic engineering. Although reduced versions of this problem have been considered, assuming certain topologies and kinetics for the feedback regulation in simplified networks [216], the more general problem remains, to the best of our knowledge, an open question. Here we propose gIOC as a general framework that includes these situations as a closed-loop inverse optimal control formulation.
@@
Note also that modeling choices and underlying assumptions have a strong impact on the resulting optimal solutions that are to be compared to observations. For example, for dynamics with linearly affine controls, $\dot{x}(t)=A(x(t),p)+B(x(t),p)u(t)$, the optimal structure is typically bang-bang for objective functionals of the type $\int_{0}^{T}\|u(\tau)\|_{1}\;\mathrm{d}\tau$, i.e., all optimal controls take values at the lower or upper bound on $u(t)$, while there are singular arcs, i.e., values in the strict interior of the lower and upper bounds, for quadratic objective functionals such as $\int_{0}^{T}\|u(\tau)\|_{2}^{2}\;\mathrm{d}\tau$. Similarly, the way dynamics and constraints are modeled influences the structure and quantitative values of independent and dependent variables. Consider the car driving problem mentioned in the introduction and illustrated in Figure 4. A simplified optimal control formulation would be: starting from rest, what is the optimal acceleration profile $a(t)$, between limits $a_{M}$ and $-a_{M}$, to go in a straight line from A to B in minimum time, stopping at B? The solution is bang-bang: maximum acceleration $a_{M}$ until the midpoint, then maximum decceleration $-a_{M}$ until the final time. This bang-bang policy is obviously not realistic because it involves abrupt switching between extreme values, which is often not feasible in real-world applications. But it is important to note that such an unrealistic bang-bang policy arises from our simplified modeling. If we consider a more refined car model, compare Figure 5, where the dynamics, path constraints, and physical limitations of actuators and the engine are considered, one obtains different optimal control policies. The resulting velocities match experimental observations quite well [217, 218], also, compare Figure 9. A similar situation is found in the linear metabolic pathway considered by Klipp et al. [55]. When simplified models are used, the solutions are bang-bang. However, when more realistic models are used, e.g., introducing kinetics for gene expression, the optimal policies may be smoother and capture the enzyme dynamics linked to the metabolic activity much better [84, 59, 58, 207].
@@ 3.1.6 A practical gIOC instance
The unknown parts of the inner level optimal control problem are modeled using the convex multipliers $w$. As discussed above, they may also depend on time in the case of switched optimality principles and are only one particular way to write down the goal to identify objective function candidates contributing to Pareto optimality and to identify activity of constraints. On the one hand the problem formulation is restrictive in the interest of a clearer presentation and might be further generalized, e.g., to multi-stage formulations involving partial differential-algebraic, delay differential, or stochastic differential equations. On the other hand, problem (8) is quite generic and allows, e.g., the consideration of switched systems, periodic processes, different underlying function spaces $\mathcal{X}$ and $\mathcal{U}$, and the usage of universal approximators such as neural networks as candidate functions. The proposed workflow is illustrated in Figure 6.
@@ 3.2 Examples of forward control problems
We consider two simple optimal control problems as examples for inner optimization problems in (8) to illustrate the effects of modeling on optimal solutions $y^{*}$.
@@
(7(d)): The solution for problem (17) modeling an external (unknown) switch at time $t=1$ and $x_{1}(1)\approx 2$. The switch from the weighted to a minimum time objective is clearly visible in the kink of the trajectory.
(7(c)): The solution for problem (18) modeling an internal (known) switch at time $t=1$ and $x_{1}(1)\approx 2.4$. Again, the switch from the weighted to a minimum time objective is clearly visible in the kink of the trajectory, but the solution differs from the solution to problem (15) also for times $t\leq 1$.
@@ 3.2.1 Brachistochrone
as well as alternative optimal control problems minimizing the integrated square angle,
@@
takes the switch of the objective function at time $1$ into account when optimizing. The optimal solutions for all six problems projected to the $(x_{1},x_{2})$ space are plotted in Figure 7. Associated optimal values $T^{*}$ are 1.44, 5.90, 1.83, 1.51, 2.59, and 8.67, respectively.
The main task of gIOC is the inverse problem: given the solutions shown in Figure 7, and assuming corresponding observations with measurement noise, is it possible to automatically and systematically derive the optimal control problem that was optimized? For the brachistochrone problem a practical analogy is somewhat unrealistic, one should rather think of an intellectual puzzle. The main take-away of this brief study is an intuition about how different objectives, constraints, and assumptions on switching the optimality principle during the observed time horizon may impact the outcome.
@@ 3.2.2 A metabolic pathway problem
In the literature, other dynamics have also been investigated [58], such as Michaelis-Menten kinetics
@@
(b) Optimal solution for problems (24), but with Michaelis-Menten kinetics (21-23).
@@
(8(a)): solution of the time minimal objective function (24) as already illustrated in Figure 2. Due to the particular bang-bang structure it is also optimal for the enzyme activiation minimizing function (25). (8(b)): modifying the assumed ODE model from (19) to (21-23) results in a structurally similar optimal control, but different differential states $x^{*}$ (note the different scale on the y-axis). (8(c)): Optimal solution for the weighted sum problem (26) with structurally different optimal controls. The reduction in $\int_{0}^{T}u^{2}_{1}(\tau)+u^{2}_{2}(\tau)+u^{2}_{3}(\tau)\;\mathrm{d}\tau$ leads also to a reduction in use of intermediate metabolites $x_{1}$ and $x_{2}$. It is balanced by an increased time $T^{*}$ to reach $x_{3}(T^{*})=0.9$. (8(d)): The differential states from (8(c)) can be arbitrarily closely approximated, although the chattering bang-bang control functions are structurally different from those in (8(c)). This shows the importance of choosing the observables for gIOC, especially considering that the variables $y^{*}$ can typically only be observed with noise.
@@ 3.3 Methods for gIOC
To our knowledge, we are the first to formulate (1) in full generality. However, algorithms for special cases have been investigated in the literature, which can serve as building blocks for a future gIOC methodology.
@@ 3.3.1 Model identification
A gIOC subproblem that is very relevant and actively researched is model identification [204]. Dynamic systems theory is a powerful tool for modeling and understanding complex systems in nature and engineering. However, developing mechanistic dynamic models of biological systems is particularly challenging, as we can not rely on first principles in the same way as in physics. Consequently, systematic model development is one of the key open problems in mathematical biology. Model discovery, i.e., the symbolic reconstruction of equations from data, is a highly desirable approach to automating the development of dynamic models. A number of different statistical and machine learning frameworks have been considered for model discovery, including symbolic regression, grammar-based methods, sparse regression, neural networks, Gaussian process regression, and Bayesian approaches, as reviewed in [222, 223, 224, 204, 225]. The sparse identification of nonlinear dynamics (SINDy) algorithm has been particularly successful [226, 227]. Other approaches solve the symbolic regression problem via mixed-integer optimization [228, 229, 230, 231]. There are many open research questions in model discovery and a significant gap in the literature concerning the concurrent identification of objective function, constraints, and dynamics. Given this ambitious goal, the availability and quality of data are essential for well-posedness of gIOC. Here, (structural and practical) identifiability, observability, and the propagation of prior error distributions towards those of quantities of interest play an important role.
@@
Identifiability and observability are crucial structural properties in the dynamic modeling of biological processes [232, 233, 234], and also play a key role in the more general problem of data-driven model discovery [235]. In standard model identification, structural identifiability refers to the ability to uniquely determine the values of the model parameters from the available input-output data, even in the absence of measurement noise. It is a fundamental requirement for reliable parameter estimation and model validation. Although methods have been proposed for handling situations with time-varying inputs [236], to the best of our knowledge there is no procedure for testing the equivalent of this property in the more general problem of inverse optimal control. However, such issues have been detected by inspection in IOC problems [157], therefore we suspect they will also be present in the gIOC framework presented here. Even if we could ensure structural identifiability, one also needs to consider practical identifiability, which in dynamic modeling refers to the ability to determine the values of model parameters effectively and reliably based on available experimental data (in gIOC the equivalent property would also involve the unique identification of controls, cost function, and constraints). The key idea here is that data needs to be rich in information, thus both the quality and quantity of the data matter. Considering gIOC, non-informative data will result in non-unique identification of the optimality principle, even if there are no structural issues.
Optimum experimental design is a systematic approach for obtaining useful data. Surveys of the design of optimal experiments in biological processes can be found in [237, 238]. Early works in cell signaling are [239, 240], and for the first time with experimental validation, [241]. The question of whether an observable should be measured at a given time or not results in another layer of mixed-integer optimal control. Compare [242] for a discussion of theoretical properties and numerical methods. Designing optimal experiments not only for parameter estimation but also for the concurrent discrimination of objectives, constraints, and models is completely open.
3.3.2 Bi-level optimization
If we focus on identifying objective functions from data and on the bi-level structure of (1), IOC and the related IRL [243] come to mind. An extensive historical review of these two approaches, tracing back to early work of Kalman on closed loop IOC for linear-quadratic systems [219], is provided in [244]. The authors also classify and discuss similarities and differences between IOC and IRL. However, they view IOC like Kalman, as related to stabilizing feedback controls, which is not our focus, as discussed in Section 3.1.5. We perceive other differences between IRL and IOC. While many RL concepts, such as those in games like chess or Go, rely on the availability of producing data via simulation, the data fitting in IOC is defined via external observational data. Therefore, data $\eta$ might not be available in abundance and can be expensive to obtain. IRL typically models dynamics as a Markov decision process or a deep neural network, requiring a large amount of data to avoid overfitting [243]. In contrast, IOC works with domain knowledge in the form of mechanistic or hybrid models using differential equations. Our interpretation of gIOC is inspired by and closely aligned with the work of [245, 71, 246, 247]. The authors formalized a bi-level formulation for the special case of fully known dynamics, switching structure, and active constraints.

### v2

@@ 3.1.3 Constraints and the feasible set $\mathcal{Y}$
While the first two types are typically addressed by a modeling expert, the third type is something that needs to be considered in the inverse optimization problem. It comprises vectors of general inequality constraints $g(y)\geq 0$. They allow to formulate lower and upper bounds on variables, but also more general relationships. Of particular importance is whether an inequality constraint $g_{i}(y)\geq 0$ is active in an optimal point, i.e., $g_{i}(y^{*})=0$, or inactive, i.e., $g_{i}(y^{*})>0$. Only active constraints have a local influence on the optimum (i.e., without this constraint the optimal point $y^{*}$ could be different). In a practical setting, if it is unclear whether the data $\eta$ was produced by an optimal, but constrained process, one could simply solve two problems – one with the constraint $g_{i}(y)\geq 0$ on the inner level, and one without. The better match to the data $\eta$ is then the more likely explanation. However, the number of possible combinations grows quickly if there are several candidate constraint functions; therefore, we model the inclusion of constraints again as a degree of freedom for the optimization on the outer level of (1). Assume that $n_{g}$ constraint function candidates $g_{i}(\cdot)$ are known, but it is unknown and of interest, if they were active and played a role in the observed process. One possibility to model them is via binary weight variables $w_{i}\in\{0,1\}$ as
@@ 3.1.5 Further considerations
Optimality principles in biology seem to be nested in a hierarchical way, as discussed above. Contributing to the main goal of maximizing fitness, biological systems may have modi operandi that are optimal for a specific function or task and the considered scale and environmental setting. Let us assume the simplified situation of a lion who knows at least three different kinds of behavior: relaxing (energy consumption minimized), sneaking up on prey (observability minimized), and hunting (time-to-target minimized). Obviously, all of them contribute to the main goal of maximizing fitness. Yet, determining the exact objective function from data may be more difficult if several switching objectives were used in the observed time horizon. Such switches can be further classified. They may either be due to external circumstances (for example, the predator suddenly observes a prey and switches from relaxing to foraging mode) or because several consecutive stages like sneaking and hunting are optimized together when foraging. Mathematically, this difference is important. In the first case, two different optimal policies are concatenated, while in the second case the optimal policy considers all stages in one go. The difference will be illustrated in Section 3.2. Additionally, the current situation (modeled via parameters $p$) has an influence. A resting lion will react differently to a prey depending on if it is hungry or not. But in both cases, the overall cost function (fitness towards survival and reproduction) is the same.
@@
In most situations where optimality principles are used to explain (or predict) dynamics, it is sufficient to consider the open loop optimal control problem, where in addition to the cost function we obtain the optimal control policies as a function of time, $u(t)$. For example, in the case of the simple linear metabolic pathway considered by Klipp et al. [55], we can obtain the time-dependent enzyme profiles and the optimality principle consistent with the data by using inverse optimal control [162]. This can be satisfactory if we simply want to explain or predict the observed dynamic behavior. However, if we want to obtain more information about the involved regulatory mechanisms, it should be noted that in reality, the enzyme concentrations depend on the expression of genes, which are themselves influenced (activated or inhibited) by the concentrations of other metabolites and enzymes, or even the expression of other genes. Thus, one might want to infer the regulatory network compatible with the observed behavior. This would be particularly useful in biotechnological and biomedical applications where we seek to change such regulations using genetic engineering. Although reduced versions of this problem have been considered, assuming certain topologies and kinetics for the feedback regulation in simplified networks [221], the more general problem remains, to the best of our knowledge, an open question. Here we propose gIOC as a general framework that includes these situations as a closed-loop inverse optimal control formulation.
@@
Note also that modeling choices and underlying assumptions have a strong impact on the resulting optimal solutions that are to be compared to observations. For example, for dynamics with linearly affine controls, $\dot{x}(t)=A(x(t),p)+B(x(t),p)u(t)$, the optimal structure is typically bang-bang for objective functionals of the type $\int_{0}^{T}\|u(\tau)\|_{1}\;\mathrm{d}\tau$, i.e., all optimal controls take values at the lower or upper bound on $u(t)$, while there are singular arcs, i.e., values in the strict interior of the lower and upper bounds, for quadratic objective functionals such as $\int_{0}^{T}\|u(\tau)\|_{2}^{2}\;\mathrm{d}\tau$. Similarly, the way dynamics and constraints are modeled influences the structure and quantitative values of independent and dependent variables. Consider the car driving problem mentioned in the introduction and illustrated in Figure 4. A simplified optimal control formulation would be: starting from rest, what is the optimal acceleration profile $a(t)$, between limits $a_{M}$ and $-a_{M}$, to go in a straight line from A to B in minimum time, stopping at B? The solution is bang-bang: maximum acceleration $a_{M}$ until the midpoint, then maximum decceleration $-a_{M}$ until the final time. This bang-bang policy is obviously not realistic because it involves abrupt switching between extreme values, which is often not feasible in real-world applications. But it is important to note that such an unrealistic bang-bang policy arises from our simplified modeling. If we consider a more refined car model, compare Figure 5, where the dynamics, path constraints, and physical limitations of actuators and the engine are considered, one obtains different optimal control policies. The resulting velocities match experimental observations quite well [222, 223], see also Figure 9. A similar situation is found in the linear metabolic pathway considered by Klipp et al. [55]. When simplified models are used, the solutions are bang-bang. However, when more realistic models are used, e.g., introducing kinetics for gene expression, the optimal policies may be smoother and capture the enzyme dynamics linked to the metabolic activity much better [84, 59, 58, 212].
@@ 3.1.6 A practical gIOC instance
The unknown parts of the inner level optimal control problem are modeled using the convex multipliers $w$. As discussed above, they may also depend on time in the case of switched optimality principles and are only one particular way to write down the goal to identify objective function candidates contributing to Pareto optimality and to identify activity of constraints.
Problem formulation (8) is given in continuous time. It is also possible to formulate a similar discrete time version. Mathematically, for the choice of constant controls $u_{k}$ and
$F(x_{k},u_{k},p):=x_{k}+\int_{t_{k}}^{t_{k+1}}f(x(\tau),u_{k}\;\mathrm{d}\tau$ (9)
the solutions of the continuous time system
$\dot{x}(t)=f(x(t),u_{k},p),\quad x(t_{k})=x_{k},\quad t\in[{t_{k},t_{k+1}}]$
and of the discrete time system
$x_{k+1}=F(x_{k},u_{k},p)$ (10)
are identical in the sense of $x(t_{k})=x_{k}$. Thus, variants of problem (8) that are based on (10) can be found in the literature. Yet, there may be practical differences and community-dependent preferences. The discrete time version may be more convenient because the time grid $t_{k}$ can be chosen matching the data $\eta$ and many inverse reinforcement learning algorithms have been formulated for discrete time problems. Discrete-time formulations can be effectively used with mechanistic models as long as the discretization preserves the essential physical meaning and time scales of the underlying continuous processes through appropriate sampling rates and numerical methods. While the implementation may be discrete, the model’s structure and parameters should still be derived from and interpreted in terms of the continuous-time physical processes they represent. Solution properties such as bang-bang behavior are better discussed in continuous time. The continuous time version is also closer to a direct mechanistic modeling and thus preferred in systems biology. We will therefore continue with the notation of (8).
On the one hand, (8) is restrictive in the interest of a clearer presentation and might be further generalized, e.g., to multi-stage formulations involving partial differential-algebraic, delay differential, or stochastic differential equations. On the other hand, problem (8) is quite generic and allows, e.g., the consideration of switched systems, periodic processes, different underlying function spaces $\mathcal{X}$ and $\mathcal{U}$, and the usage of universal approximators such as neural networks as candidate functions. The proposed workflow is illustrated in Figure 6.
@@ 3.2 Examples of forward control problems
Molecular and cellular biological systems present unique challenges and characteristics when viewed through the lens of optimal control. Unlike engineered systems, biological processes are the result of evolution rather than deliberate design. In addition, these systems often operate under multiple competing objectives. For instance, cells must balance energy efficiency against speed of response, or growth rate against robustness to environmental changes. In this context, it is worth providing an overall description of such characteristics and challenges for the three key elements of an optimal control formulation discussed in Sections 3.1.13.1.3: dynamics, objective function and constraints.
The dynamics in molecular systems biology typically involve mass-action kinetics, Michaelis-Menten enzyme kinetics, and Hill functions for describing regulatory interactions [224, 44]. These give rise to nonlinear ordinary differential equations, often with multiple time scales. A distinctive feature is that reaction rates are constrained to be non-negative, and species concentrations must remain within physiologically feasible ranges. Furthermore, many biological processes exhibit switch-like behavior, leading to dynamics that can be approximated by hybrid systems combining continuous and discrete elements.
Candidate objective functions in biological systems often reflect fundamental physical and evolutionary constraints. These may include minimization of energy expenditure (often expressed as ATP consumption), maximization of growth rate or biomass production, minimization of transition time between metabolic states, or maximization of robustness to perturbations [36, 79, 225]. In many cases, these objectives must be considered simultaneously, leading to multi-objective optimization problems [58]. The relative weights between different objectives may themselves be state-dependent, reflecting the cell’s ability to adapt its priorities based on environmental conditions.
Biological constraints are particularly complex and multifaceted [177]. At the molecular level, there are thermodynamic constraints on reaction directions and rates [93], as well as physical constraints on enzyme concentrations and metabolite levels [55, 98, 47]. Cellular machinery has limited capacity, leading to resource allocation constraints [226, 146]. For instance, ribosomes must be shared between different protein production processes. There are also temporal constraints, as many biological processes must complete within specific time windows, such as cell cycle phases. Additionally, homeostatic constraints require certain variables to be maintained within narrow ranges despite environmental fluctuations.
We focus on illustrating some fundamental concepts. For this, we start with a classical, non-biological example that is ideally suited to visualize the impact of objective function and constraints for the inner optimization problems in (8) on the optimal solution $y^{*}$ in two dimensions. As a next step, we consider a basic example from cell biology to highlight additional general issues arising in gIOC.
@@
(7(c)): The solution for problem (19) modeling an external (unknown) switch at time $t=1$ and $x_{1}(1)\approx 2$. The switch from the weighted to a minimum time objective is clearly visible in the kink of the trajectory.
(7(d)): The solution for problem (20) modeling an internal (known) switch at time $t=1$ and $x_{1}(1)\approx 2.4$. Again, the switch from the weighted to a minimum time objective is clearly visible in the kink of the trajectory, but the solution differs from the solution to problem (17) also for times $t\leq 1$.
@@ 3.2.1 Brachistochrone
as well as alternative objectives minimizing the integrated square angle,
@@
takes the switch of the objective function at time $1$ into account when optimizing. The optimal solutions for all six problems projected to the $(x_{1},x_{2})$ space are plotted in Figure 7. Associated optimal values $T^{*}$ of (1520) are 1.44, $\infty$ (if unbounded), 1.83, 1.51, 2.59, and 8.67, respectively.
The main task of gIOC is the inverse problem: given the solutions shown in Figure 7, and assuming corresponding observations with measurement noise, is it possible to automatically and systematically derive the optimal control problem that was optimized? For the brachistochrone problem a practical analogy is unrealistic, one should rather think of an intellectual puzzle. The main take-away of this brief study is an intuition about how different objectives, constraints, and assumptions on switching the optimality principle during the observed time horizon may impact the outcome.
@@ 3.2.2 A metabolic pathway problem
In the literature, also different dynamics have been investigated [58], such as Michaelis-Menten kinetics
@@
(b) Optimal solution for problems (26), but with Michaelis-Menten kinetics (23-25).
@@
(8(a)): solution of the time minimal objective function (26) as already illustrated in Figure 2. Due to the particular bang-bang structure it is also optimal for the enzyme activiation minimizing function (27). (8(b)): modifying the assumed ODE model from (21) to (23-25) results in a structurally similar optimal control, but different differential states $x^{*}$ (note the different scale on the y-axis). (8(c)): Optimal solution for the weighted sum problem (28) with structurally different optimal controls. The reduction in $\int_{0}^{T}u^{2}_{1}(\tau)+u^{2}_{2}(\tau)+u^{2}_{3}(\tau)\;\mathrm{d}\tau$ leads also to a reduction in use of intermediate metabolites $x_{1}$ and $x_{2}$. It is balanced by an increased time $T^{*}$ to reach $x_{3}(T^{*})=0.9$. (8(d)): The differential states from (8(c)) can be arbitrarily closely approximated, although the chattering bang-bang control functions are structurally different from those in (8(c)). This shows the importance of choosing the observables for gIOC, especially considering that the variables $y^{*}$ can typically only be observed with noise.
@@ 3.3 Methods for gIOC
Solving (1) in full generality is a very difficult task. However, algorithms for special cases have been investigated in the literature, which can serve as building blocks for a future gIOC methodology. We discuss the relations to system identification and interpretation in Sections 3.3.1 and 3.3.2 and to inverse reinforcement learning in Section 3.3.3, before we address the bi-level structure of (1) in 3.3.4, how to solve forward control problems in 3.3.5, and recent extensions involving inference of constraints or multi-phase objectives in 3.3.6. We discuss identifiability issues in Section 3.3.7 and illustrate the main concepts with examples in 3.3.8.
@@ 3.3.1 Model identification
A gIOC subproblem that is very relevant and actively researched is model identification (or system identification), the process of building mathematical models of dynamic systems using measured data [230, 224]. Dynamic systems theory is a powerful tool for modeling and understanding complex systems in nature and engineering. Dynamic systems theory provides the language, concepts, and framework for describing how systems evolve over time, while model identification is the practical process of finding a specific mathematical model within that framework that matches a real-world system. Model identification involves determining both the structure (the form of the equations) and parameters of a model that best explain observed input-output behavior of the system. This process typically includes selecting model candidates, designing informative experiments, collecting data, estimating parameters, and validating the resulting model against independent data [231]. Dynamic model identification distinguishes between mechanistic "white-box" models derived from first principles for physical insight, and empirical "black-box" models that are purely data-driven for flexibility at the cost of interpretability. In practice, "grey-box" models offer a pragmatic compromise, integrating known theory with data-driven components to balance predictive accuracy with scientific meaning.
By encoding the underlying physical and (bio)chemical principles, mechanistic models provide the crucial advantages of offering deep scientific insight and the power to make reliable predictions far beyond the original experimental conditions. However, developing mechanistic dynamic models of biological systems is particularly challenging, as we cannot rely on first principles in the same way as in physics. Consequently, systematic model development is one of the key open problems in mathematical biology.
@@
Model discovery, i.e., the symbolic reconstruction of equations from data, is a highly desirable approach to automating the development of dynamic models. A number of different statistical and machine learning frameworks have been considered for model discovery, including symbolic regression, grammar-based methods, sparse regression, neural networks, Gaussian process regression, and Bayesian approaches, as reviewed in [232, 233, 234, 209, 235]. The sparse identification of nonlinear dynamics (SINDy) algorithm has been particularly successful [236, 237]. Other approaches solve the symbolic regression problem via mixed-integer optimization [238, 239, 240, 241].
There are many open research questions in model discovery, including its generalization for the concurrent identification of objective function, constraints, and dynamics in inverse problems. For such ambitious goals, the availability and quality of data are essential. Here, identifiability and observability are crucial structural properties [242, 243], playing also a key role in the general problem of data-driven model discovery [244]. Identifiability and observability are two fundamental properties of dynamic models, each with a theoretical and a real-world instance. Identifiability concerns whether model parameters can be uniquely determined from data; structural identifiability is the theoretical possibility of doing so with perfect data based on the model’s equations, while practical identifiability assesses if parameters can be reliably estimated from actual, noisy experimental data [243]. Similarly, observability addresses whether a system’s internal states can be deduced from its outputs; structural observability is the theoretical possibility of this reconstruction, whereas practical observability is the ability to effectively estimate those states in the face of real-world measurement limitations [245].
Consequently, a model might be structurally identifiable but practically unidentifiable due to measurement noise, insufficient data quality or quantity, high parameter correlation, or poor experimental design. This distinction is crucial in biological systems where measurements are often sparse, noisy, and expensive to obtain. Furthermore, practical identifiability analysis can guide experimental design by revealing which measurements and perturbations are most informative for parameter estimation, and can help determine confidence intervals for parameter estimates. The interplay between these two forms of identifiability is particularly important in complex biological models where parameters often have mechanistic interpretations, and their accurate estimation is crucial for model prediction and control [242, 224].
@@
When considering the more general problem of inverse optimal control (IOC), these identifiability issues take on new dimensions. While methods have been proposed for testing structural identifiability in standard models, even with time-varying inputs [246], to the best of our knowledge there is no procedure for testing the equivalent of this property for IOC problems. Such issues have, however, been detected by inspection [162], suggesting they will also be present in generalized IOC frameworks. Moreover, even with structural identifiability ensured, practical identifiability remains a key hurdle. In the context of IOC, non-informative data will result in the non-unique identification of the entire optimality principle—including the cost function, constraints, and controls—not just the dynamic model parameters.
3.3.2 Model interpretability
Another key property is model interpretability [244, 247, 248], i.e. the ability to understand and explain a model’s structure and behavior in terms of known biological mechanisms and physical principles. An interpretable model uses mathematical expressions that correspond to real biological processes (such as enzyme kinetics, binding interactions, or regulatory relationships) rather than abstract mathematical expressions that may fit the data but lack biological meaning. This means that not only should the model’s predictions be accurate, but its structure and parameters should provide insights into the underlying biological mechanisms, allowing scientists to connect mathematical terms with specific molecular or cellular processes.
The interplay between interpretability, identifiability, and observability in biological modeling presents a complex challenge that lies at the heart of mechanistic model development [249]. A mechanistically interpretable model should reflect actual biological processes. However, the pursuit of mechanistic meaning must be balanced with the mathematical properties of identifiability and observability to ensure the model’s practical utility. While identifiability and observability are essential properties for the model to have predictive power and provide meaningful biological insights, there can be tension between these properties and mechanistic interpretability. For instance, a mechanistically meaningful model might include detailed enzyme kinetics with multiple parameters, but this complexity could lead to practical unidentifiability. Conversely, a mathematically simpler model that ensures identifiability might lack biological meaning or fail to capture important mechanistic details.
The challenge is to find representations that maintain mechanistic interpretability while ensuring identifiability and observability. This might involve careful reparametrization of mechanistic terms, lumping of parameters, or reduction of model complexity while preserving essential biological features [249]. For example, a complex Michaelis-Menten mechanism might need to be simplified to a first-order reaction under certain conditions, or multiple sequential steps might need to be represented as a single effective process. The key is to make such simplifications in ways that preserve the biological meaning of the model while ensuring its mathematical properties allow for meaningful parameter estimation and state prediction. This balance is particularly important in the context of automated model discovery methods, where the challenge is to encode both mechanistic constraints and mathematical requirements into the model selection process. Initial progress toward addressing these challenges has been reported in [244] for the case of automatic model discovery, but extending these results to the general inverse optimal control framework remains an open research direction.
3.3.3 Inverse Reinforcement Learning and IOC
Inverse Reinforcement Learning (IRL) is a machine learning technique that infers an agent’s underlying reward function by observing its behavior or demonstrations. Instead of being given a reward, IRL works backward to discover what objective the agent was optimizing to produce its observed actions [250, 251]. This general definition of inferring underlying objectives from observed behavior is very similar to the one of IOC, up to the point that both concepts are sometimes used interchangeably in the literature, e.g., “Inverse optimal control, also known as inverse reinforcement learning, …” [252, 253].
An extensive historical review of the two approaches, tracing back to early work of Kalman on closed loop IOC for linear-quadratic systems [227], is provided in [254]. The authors also classify and discuss similarities and differences between IOC and IRL. However, they view IOC like Kalman, as mainly related to stabilizing feedback controls, which is not our focus, as discussed in Section 3.1.5.
In our view, and closer to another IRL survey [255], the main differences are linked to terminology and historical context (IOC has roots in classical control theory, where systems are typically deterministic and continuous-time; IRL emerged more from the machine learning community), underlying system models (IOC traditionally assumes a state-space model; IRL explicitly assumes a Markov Decision Process (MDP), accounting for stochastic environments and agent experience), and computational approaches (IOC is based on optimization and control theory, IRL often applies stochastic optimization or machine learning algorithms). An additional practical difference arises from the availability of data. Modeling with MDP or deep neural networks in IRL requires a large amount of data to avoid overfitting [250]. Many RL concepts, such as those in games like chess or Go, rely on the availability of producing data via simulation. In contrast, in molecular biological applications, data is often sparse and possibly expensive to obtain. Also, interpretability of results may be more of an issue when compared to other fields of application, such as robotics or games. Therefore mechanistic or hybrid models using differential equations may be better suited in this context.
Notwithstanding, we share the impression that “the fields of IRL and IOC have essentially merged and become interchangeable” [255], if general concepts and algorithmic ideas are concerned. This convergence is driven by model identification as outlined in Section 3.3.1 and by a trend towards “hybrid modeling”: mechanistic ordinary differential equation (ODE) models are often enhanced with data-driven surrogate submodels (UDE) on the one hand [213], and physics-informed universal approximators (PINN) [210] have become an important field of research on the other hand. This convergence of prediction models is stimulating similar algorithmic developments for both IOC and IRL.
Converging modelsAlgorithmsODEUDEPINNMDPIOCgIOCIRL
So, in theory and by making use of (9) to represent continuous time dynamics as a special case of a MDP, all IRL algorithms can also find application in IOC. IRL methods, such as maximum entropy [256] or GAIL, have been surveyed in [255]. It is an open and fascinating research question for which types of underlying prediction models these algorithms can be applied successfully in practice.
We believe that the unique demands of systems biology (such as the need for interpretable mechanistic differential equation models, the vastly different scale and quality of biological training data, and the availability of efficient nonlinear optimal control solvers) necessitate a specialized approach to gIOC algorithm development. This effort must extend beyond a straightforward adaptation of IRL methods.
3.3.4 Algorithmic approaches to the bi-level problem
Our interpretation of gIOC is inspired by and closely aligned with the work of [257, 71, 258, 259]. The authors formalized a bi-level formulation for the special case of fully known dynamics, switching structure, and active constraints.